Environmental Consulting can buy some field equipment for $48,000 or lease it for 2 years at $28,000 per year. Salvage value of the equipment after 2 years would be $7,000. What is the IRR of the leasing option?

( Present your answer in percentage form with 2 decimal places and without the % sign)

economics question

be clear, show steps, solve in 30 minutes for thumbs up rating
Answer #1

Solution :- For the calculation of IRR, Take two trial discount rates. One discount rate at which the present value of cash inflows are greater than the present value of cash outflow and another discount rate at which the present value of cash inflows are lesser than the present value of cash outflow..

At 15 % discount rate :-

Present value of cash inflows = 28000 / (1 + 0.15)1 + 28000 / (1 + 0.15)2 + 7000 / (1 + 0.15)2

= 28000 / (1.15)1 + 28000 / (1.15)2 + 7000 / (1.15)2

= 28000 / 1.15 + 28000 / 1.3225 + 7000 / 1.3225

= 24347.83 + 21172.02 + 5293.01

= $ 50812.85 (Rounded off to $ 50813)

Present value of cash outflow = $ 48,000. (Given in the question)

At 20 % discount rate :-

Present value of cash inflows = 28000 / (1 + 0.20)1 + 28000 / (1 + 0.20)2 + 7000 / (1 + 0.20)2

= 28000 / (1.20)1 + 28000 / (1.20)2 + 7000 / (1.20)2

= 28000 / 1.20 + 28000 / 1.44 + 7000 / 1.44

= 23333.33 + 19444.44 + 4861.11

= $ 47638.88 (Rounded off to $ 47639)

Present value of cash outflow = $ 48,000.(Given in the question)

By inter-polation, Internal rate of return (IRR) calculated as follows :-

= 15 % + (20 % - 15 %) * (50813 - 48000) / (50813 - 47639)

= 15 % + 5 % * 2813 / 3174

= 15 % + 5 % * 0.8863 (approx)

= 15 % + 4.43 % (approx)

= 19.43 %

Conclusion :- IRR of leasing decision (in percent) = 19.43